Reports 1-1 of 1 Clear search Modify search
Infrastructures (Air Conditioning)
soldani - 15:08 Wednesday 28 July 2021 (52685) Print this report
CEB AHU outlet fan new frequency
To try to compensate the 1 Pa drop in outlet-air pressure from the AHU in the Central Hall at 2.30 pm (L.T.) the new frequency of 46 Hz was set (+1 Hz compared to the previous one).
The delivery pressure went from 119 to 126 Pa.
Return fan frequency remained unchanged at 45 Hz.

Images attached to this report
Search Help